OpenStudy (anonymous):

The lengths of two sides of a triangle are shown below: Side 1: 5x2 + 2x - 3 Side 2: 8x - 3x2 + 2 The perimeter of the triangle is 2x3 - 5x2 + 4x - 8. Part A: What is the total length of the two sides, 1 and 2, of the triangle? (4 points) Part B: What is the length of the third side of the triangle? (4 points) Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer. (2 points)

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@perl

OpenStudy (ilovecake):

I am so sorry but, I don't know how to do this. sorry... :(

OpenStudy (radar):

Part A. Just add side 1 and side 2, combining similar terms and being careful of signs. Side 1: 5x2 + 2x - 3 + Side 2: 8x - 3x2 + 2 =----------------

OpenStudy (anonymous):

2x2+10x-1

OpenStudy (radar):

Yes, now subtract that sum from the perimeter and you will have Part B.

OpenStudy (anonymous):

so wait thats part A?

OpenStudy (anonymous):

Part B: 2x^3-7x^2+14x-16?

OpenStudy (radar):

I did not get that You are subtracting 2x2+10x-1) from 2x3 - 5x2 + 4x -8 (the perimeter. I differ with you on the x term and the numeric (the 14x -16)

OpenStudy (anonymous):

4x+16?

OpenStudy (anonymous):

no just subtracting so 4x

OpenStudy (anonymous):

or -6x

OpenStudy (anonymous):

okay so its 2x^3-3x^2-6x

OpenStudy (radar):

4x -10x = -6x

OpenStudy (radar):

-8 -(-1) = -7

OpenStudy (anonymous):

so 2x^3-3x^2-6x-7?

OpenStudy (radar):

All together 2x3 - 7x2 -6x - 7

OpenStudy (anonymous):

okay okay i understand i got it now. what about c?

OpenStudy (radar):

Yes, demonstrated by the fact when you combine the side 1, side 2 and the result of side 3 you get the given perimeter.

OpenStudy (anonymous):

im confused how is that part C?

OpenStudy (radar):

That is part C.

OpenStudy (anonymous):

okay

OpenStudy (radar):

You could show this by adding side 1 , side 2 and the answer you got for Part B and showing that result is equal to the given Perimeter.

OpenStudy (anonymous):

okay your a huge help do you think you can help me finish out my study guide?

OpenStudy (anonymous):

it only has 4 more questions that im stuck on.

OpenStudy (radar):

I gotta run now as my mothers therapist is here and I need to observe.